OpenStudy (anonymous):

What is the simplest form of the expression below? f^2-25/5 * f/f-5

OpenStudy (anonymous):

f2−255ff−5 =f3−10ff =f2−10

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

im pretty sure it is

OpenStudy (anonymous):

it has to be put in similest form

OpenStudy (mathmate):

Your are missing parentheses if you mean \(\large \frac{f^2-25}{5}\times\frac{f}{f-5}\) which should have been written (f^2-25)/5 * f/(f-5) according to mathematical conventions (PEMDAS)

OpenStudy (mathmate):

hint: factor and cancel common factors. make sure to insert condition that the cancelled factor does not equal zero.